Using Maxwell’s equations in free space, derive the wave equation for the

y-component of the electric field vector.


1
Expert's answer
2021-08-29T16:55:15-0400

Maxwell’s equations in free space are given by

"\nabla\cdot {\bf E}=0\\
\nabla\cdot {\bf B}=0""\nabla\times {\bf E}=-\frac{1}{c}\frac{\partial {\bf B}}{\partial t}""\nabla\times {\bf B}=\frac{1}{c}\frac{\partial {\bf E}}{\partial t}"

The last two equations give

"\nabla\times \frac{\partial{\bf B}}{\partial t}=\frac{1}{c}\frac{\partial^2 {\bf E}}{\partial t^2}"

"-c\nabla\times \nabla\times {\bf E}=\frac{1}{c}\frac{\partial^2 {\bf E}}{\partial t^2}"

or


"\nabla\times \nabla\times {\bf E}=-\frac{1}{c^2}\frac{\partial^2 {\bf E}}{\partial t^2}"

Using identity

"\nabla\times \nabla\times {\bf E}=\nabla(\nabla\cdot {\bf E})-\nabla^2{\bf E}"

and first Maxwell’s equation, we obtain

"\nabla\times \nabla\times {\bf E}=-\nabla^2{\bf E}=-\frac{1}{c^2}\frac{\partial^2 {\bf E}}{\partial t^2}"

Finally

"\nabla^2{\bf E}-\frac{1}{c^2}\frac{\partial^2 {\bf E}}{\partial t^2}=0"

Also we have

"\nabla^2E_y-\frac{1}{c^2}\frac{\partial^2 E_y}{\partial t^2}=0"